OpenStudy (anonymous):

Please help! What are the upper and lower bounds for the function: f(x) = sin x+ 2? Are they asking for the local max and local min?

OpenStudy (tyteen4a03):

Remember that sin() has an upper value of 1 and a lower value of -1.

OpenStudy (anonymous):

The answers can't be 3 (upper value) and 1 (lower value)... can they? That seems way too easy.

OpenStudy (tyteen4a03):

@pathosdebater Correct. It's that easy.

OpenStudy (anonymous):

But when I type the equation into a graphing calc, I get the local min = -1 and the local max = 1

OpenStudy (anonymous):

Should the equation in the calculator be y = sin(x+2) or y = sin(x) + 2?

OpenStudy (anonymous):

I got 3 as the local max and 1 as the local min when I put y = sin(x) + 2

OpenStudy (tyteen4a03):

It sounds like the graphic calculator is lying.

OpenStudy (anonymous):

About which one?